supie:

A lawn-mowing company is trying to grow its business. It had 18 clients when they started its business and wants to increase by 4 new clients each week. Use an arithmetic sequence to write a function to represent this real-world situation and determine the range of the function for the first four weeks of data. f(x) = 4x + 14; 0 ≤ y ≤ 4 f(x) = 4x + 18; 0 ≤ y ≤ 4 f(x) = 4x + 18; 18 ≤ y ≤ 30 f(x) = 4x + 14; 18 ≤ y ≤ 30

Imagine:

So already you can mark A,D out of the question because there is 18 new clients, not 14.

Imagine:

And you are trying to gain each clients through the week, so it would be f(x) = 4x + 18; 18 ≤ y ≤ 30

TWALKER16:

f(x) =4x+18; 18≤y≤30 is correct

umm:

First week, number of clients = 30 Second week, number of clients = 30 + 6(1) Third week, number of clients = 30 + 6(2) Third week, number of clients = 30 + 6(3) Fourth week, number of clients = 30 + 6(4) Therefore\[f(x)= 6 x + 30; 0 ≤ y ≤ 4\]

supie:

Ok thanks guys!

Imagine:

No problem!